OpenStudy (anonymous):

plseeeeeeeeee i need it urgentlyyyyyyyyyyyyyy need it very badly...very urgent

Parth (parthkohli):

Can't do integrals; sorry.

OpenStudy (anonymous):

who can do it...plse tell me

Parth (parthkohli):

UnkleRhaukus is the best in the business I think.

OpenStudy (unklerhaukus):

i really dont think i can help here

terenzreignz (terenzreignz):

An indefinite integral? Why don't you just differentiate the right-hand side and hope for the best? :D

terenzreignz (terenzreignz):

doodling... \[\large 1 + \tan(x)\tan(x+\theta) = 1+\tan(x)\left[\frac{\tan(x)+\tan(\theta)}{1-\tan(x)\tan(\theta)}\right]\]

terenzreignz (terenzreignz):

Just to spare myself from confusion, sort of, I'm letting \[\large p = \tan(\theta)\] \[\large 1 + \frac{\tan^2(x) + p\tan(x)}{1-p\tan(x)}\]

terenzreignz (terenzreignz):

\[\large =\frac{1-p\tan(x)+\tan^2(x)+p\tan(x)}{1-p\tan(x)}\]

terenzreignz (terenzreignz):

\[\large \frac{1+\tan^2(x)}{1-p\tan(x)}\]\[\large \frac{\sec^2(x)}{1-p\tan(x)}\]

terenzreignz (terenzreignz):

This looks much more palatable, don't you think? @msingh @ParthKohli @UnkleRhaukus :D

OpenStudy (anonymous):

i think i got it

terenzreignz (terenzreignz):

I was just getting to the fun part >.> Oh well, good on you :D

OpenStudy (agent0smith):

From where @terenzreignz left it, the integration should be easy; let u=1-ptanx and doo-wop that thing.

terenzreignz (terenzreignz):

doo-wop :D

OpenStudy (agent0smith):

That thing, that thing, that thiiiiiiing. Like Lauryn Hill.